Suppose the mean income of firms in the industry for a year is 85 million dollars with a standard deviation of 9 million dollars. If incomes for the industry are distributed normally:

a. What is the probability that a randomly selected firm will earn less than 96 million dollars? Round your answer to four decimal places.

b. Indicate (shade) on the curve where the asking probability would be.
Answer #1

let the income tmeans is random variable 9 X ~N(85, 92 we know that it xn N(M02) then. 2=X-M ~ NO.1 a P (X<96) = P(X-85. < 96